What will be the final temperature if a 4 x 10⁷ J of heat is transferred to a 1 kg block of aluminum initially at 20°C

1
Expert's answer
2022-02-22T10:30:18-0500

Given:

"Q=4*10^7\:\rm J"

"m=1\:\rm kg"

"t_i=20\:\rm {^\circ C}"

"c=890\:\rm J/kg\cdot K"


The amount of heat

"Q=cm\Delta t"

Hence, the final temperature of a block

"t_f=t_i+Q/cm\\
=20+4*10^7/(890*1)=44964 \:\rm {^\circ C}"

The obtained result is too large. Please check your given.